# On the slow phase for fixed-energy Activated Random Walks

### v1

@@ 1 Introduction
Activated Random Walk (ARW) is an Abelian model of interacting particles that can be described as follows. On the one-dimensional ring $\mathbb{Z}_{N}:=\mathbb{Z}/N\mathbb{Z}$, particles are distributed with initial density $\zeta$. There are two types of particles, active and sleeping particles. Active particles perform independent continuous-time random walks at rate $1$, and fall asleep at rate $\lambda$. Sleeping particles do not move, and continue to sleep until an active particle reaches the same site, and then become active again. See [14] for an introduction and main results, and [1, 2, 4, 5, 6, 7, 8, 9, 11, 12, 13, 15] for more recent developments.
As the ring $\mathbb{Z}_{N}$ is finite, it is immediate that there is fixation a.s. if, and only if, $\zeta\leq 1$. Thus, we are interested in how much time time the system will take to stabilize. Our main result is the following theorem.
Theorem 1.1.
@@
While the authors worked on this problem, two papers came out. In [1], Theorem 1.1 is proved for the $d$-dimensional torus, $d\geq 2$, using auxiliary results and arguments of [7]. Their proof uses a hierarchical argument. Fixed a “dormitory” set where the particles will eventually stabilize, the model is reduced to a model of density $1$ in the dormitory, which is then divided in a series of growing “clusters”. Stabilization of each of those clusters will affect the configuration in the other clusters, leading to a long stabilization time with high probability, even after summing over all the possible dormitory sets. A modification on this argument shows that the result is still true in the one-dimensional case. In [9] it is shown that the critical density for the fixed-energy model on $\mathbb{Z}_{N}$ and $\mathbb{Z}$ are the same. Their proof is based on the cyclical argument of [3], using a “percolation in layers” approach to show that, in every cycle, there will remain a fraction of active particles that will be sufficient to wake up the sleeping particles.
Our proof uses a different approach which is self-contained and shorter. The main strategy is to build a “carpet” where active particles can perform long excursions. Unlike [10], we do not assume a convenient initial configuration, instead our carpet is built on the fly while the system dynamics develop. We then divide the ring in large blocks, and prove that with high probability the system reaches certain configurations having many active particles in a certain region of the ring. In contrast to [3], which handles the low $\lambda$ regime, at the end of each cycle we are left inconvenient configurations connecting the source and sink regions, and need to make do with them.
This note is structured as follows. In Section 2 we give a formal description of the ARW model. In Section 3 we describe a dynamic procedure to move particles on the ring according to the ARW rules. Theorem 1.1 is proved in Section 4, using the cycle argument. In Section 5 we prove a bound on the exponential moment of the number of “bad” blocks. In Section 6 we prove a one-block estimate and in Section 7 we use bounds on the tail of excursions of random walks to estimate the drift of the sites where there will be no particles.
@@ 2 Description of the Model
$[\mathfrak{t}_{x,y}\eta](z)=\begin{cases}\eta(z)+1,&\text{if }z=y;\\ \eta(z)-1,&\text{if }z=x\\ \eta(z),&\text{otherwise,}\end{cases}\qquad[\mathfrak{t}_{x,\mathfrak{s}}\eta](z)=\begin{cases}\mathfrak{s},&\text{if }z=x\text{ and }\eta(x)=1;\\ \eta(z),&\text{otherwise.}\end{cases}$
@@
The Abelian property of the ARW (Lemma 2.4 of [14]) states that if $\alpha$ and $\beta$ are two legal sequences of topplings that stabilize a configuration $\eta$, then $m_{\alpha}=m_{\beta}$. In other words, the order in which legal topplings are performed does not affect the final configuration obtained.
@@ 3.1 Preliminary steps
Without loss of generality, we can assume that the initial configuration has at most only particle per site. Otherwise, because $\zeta<1$, we can a.s. topple sites with more than one particle until there are no sites with more than one particle.
Let $a\in 2\mathbb{N}$ be fixed, its value will be determined later, and let $K:=a^{2}$. For simplicity, we assume $N=(n+2)K$ for some even integer $n$. The ring $\mathbb{Z}_{N}$ will be identified with the interval $\{-K/2,-K/2+1,\dots,N-K/2-1\}$ in the obvious way. For $10\leq i\leq n+$ 1, we call the set $[iK-K/2,iK+K/2)$ the block $i$.
@@
    If there are defective sites in $[iK-K/2,iK+K/2)$, the hole is at $iK$.
@@ 3.2 Attempted emission
Suppose we choose the hot particle at block $i$. A successful emission (or just emission) occurs either when the hot particle reaches $(i+1)K$ (resp. $(i-1)K+a$) and there are no defects in block $i+1$ (resp. $i-1$); or when the hot particle reaches a vacant site in block $i+1$ (resp. $i-1$) and there are defects in block $i+1$ (resp. $i-1$). We remember that, by (P2), vacant sites are either holes or defects. In either case, block $i$ is called the emitting block and block $i+1$ (resp $i-1$) the receiving block. When an emission occurs, we say that the hot particle is emitted.
Let us describe the procedure for an attempted emission. First, we will choose the hot particle. The hot particle is chosen using the following criterion. At the beginning of an attempted emission, pick the smaller $i\in\{1,2,\dots n\}$ such that there are no defects and there is at least one thawed free particle in block $i$. If there is no such $i$, we declare the procedure finished. After $i$ is chosen, we pick one of the free thawed particles of block $i$ to be the hot particle. Note that these free thawed particles must be in $iK$ or $iK+a$ by (P6). If there is a free thawed particle at $iK$, we choose the hot particle at $iK$; else, we choose the hot particle at $iK+a$. Note that the hot particle is never chosen in blocks 00 and $n+1$.
@@
There is no frozen particle at site $iK+a$: Topple the hot particle until it is is emitted or reaches the hole in $[iK,iK+a]$. If it is emitted, we finish the attempted emission. If the hot particle reaches the hole and sleeps, we turn it into a carpet particle, move the hole to the site immediately to the right, turn the carpet particle in the new location of the hole into a free particle, and choose this particle to be hot ((P4) ensures that the new hot particle is active). If the hole reaches $iK+a$, we stop the attempted emission and declare the free particle at $iK+a$ as frozen. If this happens, we say that the attempted emission ends in a failure and, if possible a new hot particle is chosen in the way described above.
@@
-The particle makes an excursion and returns to the hole. In this case, we move the hole to the leftmost site visited by the hot particle in this excursion, turn the hot particle into a carpet one, and turn the particle at the new position of the hole into a free one. (Note that, if the hot particle made an excursion to the right of the hole, the leftmost site visited is the hole itself.)
There is a frozen free particle at $iK+a$: In this case, by the combination of (P2)(P3) and (P8), there are no defects on $[iK-K/2,iK+K/2)$, and all sites in $[iK,iK+a)$ have a carpet particle. We topple the hot particle until it is emitted. If the particle visits every site in $[iK,iK+a]$ before being emitted, we move the hole to $iK$, turn the frozen particle at $iK+a$ into a carpet particle and turn the carpet particle at $iK$ into a free particle. This finishes the attempted emission.
See Figure 1 to an example of a piece of the ring after some attempted emissions have been made.
Figure 1: Example of possible state of a piece of the ring when an attempted emission is about to begin. Filled circles are active carpet particles, empty circles are sleeping carpet particles. The squares represent free particles, one of which is frozen. The holes are indicated by underlining the site.
@@ 4 Alternating modes
We will now prove Theorem 1.1. To do so, we apply cyclically the procedure of Section 3 in two modes, A and B. At each mode, blocks 00 and $n+1$ are called sinks and blocks $n/2$ and $n/2+1$ are called sources. At the end of each mode we change the enumeration of the blocks in a way such that the sinks of mode A are the sources of mode B and vice-versa. Note that the hot particle is never chosen in the sink regions.
Let $D$ the number of defects at the end of a mode. We will define the following Condition and use it in the proof of Theorem 1.1.
Condition 4.1.
@@
Proposition 4.2.
There is $c>0$ depending on $\lambda$ such that, if at the beginning of a mode Condition 4.1 is satisfied, then at the beginning of the next mode Condition 4.1 will also be satisfied with probability at least $1-e^{-cn}$, for all $n$ sufficiently large.
We will prove Proposition 4.2 at the end of this section. Now we are ready to prove Theorem 1.1.
Proof of Theorem 1.1.
We take $11-\frac{1}{4K}<\zeta<$ 1. Therefore, at the initial configuration there are at most $n/4$ defects, and at least $43n/$ 4 blocks with one free particle. Thus, there will be at least one block without defects and with one free thawed particle, and we can start the toppling procedure.
We will call a mode successful if Condition 4.1 is met at the end of the mode. We first note that, if a mode is successful, at least one jump is made.
Using Proposition 4.2,
@@
Let $E_{n}$ be the set of blocks that realize at least one successful emission during a mode, $F(E_{n})$ be the number of frozen particles in $E_{n}$ and $F(\mathbb{Z}_{N})$ the total number of frozen particles in the ring $\mathbb{Z}_{N}$.
To prove the Proposition 4.2 we will make use of the following statements.
Proposition 4.3.
There exists a constant $c$ depending on $\lambda$ such that, if Condition 4.1 is satisfied then $\mathbb{P}(F(E_{n})\geq\frac{n}{8})\leq e^{-cn}$ for all $n$ sufficiently large.
Proposition 4.4.
There exists a constant $c>0$, depending on $\lambda$ such that, if at the beginning of the mode all the holes are at positions $iK$, for $i=1,\dots,n$, then $\mathbb{P}(F(\mathbb{Z}_{N})\geq\frac{n}{8})\leq e^{-cn}$ for all $n$ sufficiently large.
Proof of Proposition 4.2..
Note that the number of free particles minus the number of defects is conserved by the toppling procedure, therefore at the end of the mode there are still $\frac{7}{8}n+D$ free particles.
In the first mode all the holes are at positions $iK$, for $i=1,\dots,n$, hence by Proposition 4.4, the probability of having at most $n/8$ frozen particles in total at the ring is at least $1-e^{-cn}$. In this event, Condition 4.1 is met.
Now, let us show that, on the event $\{F(E_{n})\leq n/8\}$, Condition 4.1 is met at the end of the mode.
We first argue that in this event, there will be free particles inside the sink blocks. From the second mode onward, all free thawed particles begin the mode in the source blocks. Therefore, $E_{n}$ is a connected set surrounding the source, and frozen particles outside $E_{n}$ must be in the two blocks adjacent to $E_{n}$.
If $\{F(E_{n})\leq n/8\}$ happens, at the end of the mode, there are at most 2 new frozen particles outside $E_{n}$ and $85n/$ 8 particles that were frozen at the beginning of the mode. Since $85n/8+n/8+2<7n/$ 8 (if $n>16$), and the three options listed earlier are the only places where free frozen particles may be; there must have been free particles that are not frozen by the end of the mode, and this can only be achieved if free particles are emitted to a sink block.
Therefore, we conclude that, in fact, $E_{n}$ must contain the source block and be adjacent to a sink block, and since $E_{n}$ is a connected collection of blocks, the number of blocks in the complement of the union of $E_{n}$ and the sink blocks is at most $n/2$. In particular, by (P7), there are at most $n/2$ frozen particles outside $E_{n}$, hence at most $85n/$ 8 frozen particles in total.
By Proposition 4.3, $\mathbb{P}(F(E_{n})<n/8)\geq 1-e^{-cn}$, and Proposition 4.2 is proven. ∎

### v2

@@ 1 Introduction
Activated Random Walk (ARW) is an Abelian model of interacting particles that can be described as follows. On the one-dimensional ring $\mathbb{Z}_{N}:=\mathbb{Z}/N\mathbb{Z}$, we take a deterministic number $\zeta N$ of particles in arbitrary initial positions. There are two types of particles, active and sleeping particles. Active particles perform independent continuous-time random walks at rate $1$, and fall asleep at rate $\lambda$. Sleeping particles do not move, and continue to sleep until an active particle reaches the same site, and then become active again. See [14] for an introduction and main results, and [1, 2, 4, 5, 7, 6, 8, 9, 11, 12, 13, 15] for more recent developments.
As the ring $\mathbb{Z}_{N}$ is finite, it is immediate that there is fixation a.s. if, and only if, $\zeta\leq 1$. Thus, we are interested in how much time the system will take to stabilize. Our main result is the following theorem.
Theorem 1.
@@
While the authors worked on this problem, two papers came out. In [1], Theorem 1 is proved for the $d$-dimensional torus, $d\geq 2$, using auxiliary results and arguments of [6]. Their proof uses a hierarchical argument. Fixed a “dormitory” set where the particles will eventually stabilize, the model is reduced to a model of density $1$ in the dormitory, which is then divided in a series of growing “clusters”. Stabilization of each of those clusters will affect the configuration in the other clusters, leading to a long stabilization time with high probability, even after summing over all the possible dormitory sets. A modification on this argument shows that the result is still true in the one-dimensional case. In [9] it is shown that the critical density for the fixed-energy model on $\mathbb{Z}_{N}$ and $\mathbb{Z}$ are the same. Their proof is based on the cyclical argument of [3], using a “percolation in layers” approach to show that, in every cycle, there will remain a fraction of active particles that will be sufficient to wake up the sleeping particles.
Our proof uses a different approach which is self-contained and shorter. The main strategy is to build a “carpet” where active particles can perform long excursions. Unlike [10], we do not assume a convenient initial configuration, instead our carpet is built on the fly while the system dynamics develop. We then divide the ring in large blocks, and prove that with high probability the system reaches certain configurations having many active particles in a certain region of the ring. The core of the argument is that, if we take the density large enough (roughly $1-C_{1}\exp(C_{2}\lambda)$, with $C_{1}$ and $C_{2}$ constants), we can overcome the large sleeping rate by using the fact that, when a particle does not sleep, it will make a long excursion and wake up many particles. In contrast to [3], which handles the low $\lambda$ regime, at the end of each cycle we are left with inconvenient configurations connecting the source and sink regions, and need to make do with them.
This note is structured as follows. In Section 2 we give a formal description of the ARW model. In Section 3 we describe a dynamical procedure to move particles on the ring according to the ARW rules. This procedure is constructed in such a way that certain important properties are preserved, as shown in Section 4. Theorem 1 is proved in Section 5, using the cycle argument. In Section 6 we prove a bound on the exponential moment of the number of “bad” blocks. In Section 7 we prove a one-block estimate and in Section 8 we use bounds on the tail of excursions of random walks to estimate the drift of the sites where there will be no particles.
@@ 2 Description of the Model
[…]
@@
The Abelian property of the ARW (Lemma 2.4 of [14]) states that if $\alpha$ and $\beta$ are two legal sequences of topplings that stabilize a configuration $\eta$, then $m_{\alpha}=m_{\beta}$. Furthermore, the order in which legal topplings are performed does not affect the final configuration obtained.
@@ 3.1 Preliminary steps
Without loss of generality, we can assume that the initial configuration has at most one particle per site. Otherwise, because $\zeta<1$, we can a.s. topple sites with more than one particle until there are no sites with more than one particle.
Let $a\in 2\mathbb{N}$ be fixed, its value will be determined later, and let $K:=a^{2}$. For simplicity, we assume $N=(n+2)K$ for some even integer $n$. The ring $\mathbb{Z}_{N}$ will be identified with the interval $\{-K/2,-K/2+1,\dots,N-K/2-1\}$ in the obvious way. For $0\leq i\leq n+1$, we call the set [iKK/2,iK+K/2)[iK-K/2,iK+K/2) the block $i$.
@@
The density $\zeta$ will be chosen later in such a way that on average, each block contains less than one vacant site (a hole or a defect).
@@
    If there are defective sites in [iKK/2,iK+K/2)[iK-K/2,iK+K/2), the hole is at $iK$.
@@ 3.2 Attempted emission
Suppose we choose the hot particle at block $i\in\{1,\dots,n\}$. Note that the hot particle is never chosen in blocks 0 and $n+1$. A successful emission (or just emission) occurs either when the hot particle reaches $(i+1)K$ (resp. (i1)K+a(i-1)K+a) and there are no defects in block $i+1$ (resp. i1i-1); or when there are defects in block $i+1$ (resp. i1i-1) and the hot particle reaches a vacant site in block $i+1$ (resp. i1i-1). We remember that, by (P2), vacant sites are either holes or defects. In either case, block $i$ is called the emitting block and block $i+1$ (resp i1i-1) the receiving block. When an emission occurs, we say that the hot particle is emitted.
Note in particular that the hot particle may be emitted when reaching the hole of block i1i-1 (resp. $i+1$) if this hole is vacant and if there are defects in [(i1)KK/2,(i1)K][(i-1)K-K/2,(i-1)K] (resp. $[(i+1)K,(i+1)K+K/2]$).
Let us describe the procedure for an attempted emission. First, we will choose the hot particle. The hot particle is chosen using the following criterion. At the beginning of an attempted emission, pick the smaller $i\in\{1,2,\dots n\}$ such that there are no defects and there is at least one thawed free particle in block $i$. If there is no such $i$, we declare the procedure finished. After $i$ is chosen, we pick one of the free thawed particles of block $i$ to be the hot particle. Note that these free thawed particles must be in $iK$ or $iK+a$ by (P6). If there is a free thawed particle at $iK$, we choose the hot particle at $iK$; else, we choose the hot particle at $iK+a$.
@@
There is no frozen particle at site $iK+a$: Topple the hot particle until it is is emitted or reaches the hole in $[iK,iK+a]$. If it is emitted, we finish the attempted emission. If the emission occurs because the hot particle fixed a defect, we turn the hot particle into a carpet one. Otherwise, it remains free and thawed. Note that an emission may occur if the hot particle reaches a vacant hole and there are still defects in the neighbor block. If the hot particle reaches the hole and sleeps, we turn it into a carpet particle, move the hole to the site immediately to the right, turn the carpet particle in the new location of the hole into a free thawed particle, and choose this particle to be hot ((P4) ensures that the new hot particle is active). If the hole reaches $iK+a$, we stop the attempted emission and declare the free particle at $iK+a$ as frozen. If this happens, we say that the attempted emission ends in a failure and, if possible a new hot particle is chosen in the way described above.
@@
-The particle makes an excursion and returns to the hole. In this case, we move the hole to the leftmost site of $[iK,iK+a]$ visited by the hot particle in this excursion, turn the hot particle into a carpet one, and turn the particle at the new position of the hole into a free thawed one and turn this particle in the hot particle. (Note that, if the hot particle made an excursion to the right of the hole, the leftmost site visited is the hole itself.)
There is a frozen free particle at $iK+a$: In this case, by the combination of (P2)(P3) and (P8), there are no defects on [iKK/2,iK+K/2)[iK-K/2,iK+K/2), and all sites in $[iK,iK+a)$ have a carpet particle. We topple the hot particle until it is emitted. If the particle visits every site in $[iK,iK+a]$ before being emitted, we move the hole to $iK$, turn the frozen particle at $iK+a$ into a carpet particle and turn the carpet particle at $iK$ into a free thawed particle. This finishes the attempted emission.
See Figure 1 for an example of a piece of the ring after some attempted emissions have been made.
Figure 1: Example of possible state of a piece of the ring when an attempted emission is about to begin. Filled circles are active carpet particles, empty circles are sleeping carpet particles. The squares represent free particles, one of which is frozen. The holes are indicated by underlining the site. There is a frozen particle at $iK+a$.
4 Properties preserved
In this section, we prove that the properties (P1) to (P9) of a configuration are preserved by the procedure described in Section 3.
Proposition 1.
At the end of each attempted emission, each one of properties (P1) to (P9) is preserved.
Proof.
This follows from the following reasons:
(P1) We never move the hole outside $[iK,iK+a]$, and every time we move the hole to a new location, the old location ceases to be a hole. Thus, there is only one hole at each block.
(P2) This happens because sites that are either defects or holes only cease to be defects and holes if a particle reaches them and is turned into a carpet particle; and with exception of the hole, we only topple sites with at least two particles.
(P3) We only move the hole after all defects in [iKK/2,iK+K/2)[iK-K/2,iK+K/2) are fixed, because there are no defects in the block which the hot particle is chosen from.
(P4) If a particle fixes a defect, we relabel it as a carpet one and it remains active. If a particle sleeps, we move the hole to the right. We only move the hole to the left if all the particles between its new location and $iK+a$ are active.
(P5) Every time a free particle sleeps, we relabel it as a carpet particle.
(P6) We only move the hot particle, and every time the hot particle fixes a defect or sleeps, we relabel the hot particle as a carpet one. We only stop to move the hot particle when it is emitted or frozen. Thus, every time we choose a new hot particle, the old one is either relabeled as a carpet particle (if it fixed a defect), is frozen (at $iK+a$) or reached block i1i-1 or $i+1$.
(P7) If there is a frozen free particle at the block where the hot particle is chosen, the hot particle is always emitted or fixes a defect. In either case, we do not freeze any other particle.
(P8) This is because a particle is frozen if and only if the hole reaches $iK+a$, and is turned from frozen to carpet when the hole leaves $iK+a$.
(P9) We choose a free and thawed particle to be hot. When a hot particle is relabeled as a carpet one or declared frozen, a new hot particle is chosen.
@@ 5 Alternating modes
We will now prove Theorem 1. To do so, we apply cyclically the procedure of Section 3 in two modes, A and B. At each mode, blocks 0 and $n+1$ are called sinks and blocks $n/2$ and $n/2+1$ are called sources. At the end of each mode we change the enumeration of the blocks in a way such that the sinks of mode A are the sources of mode B and vice-versa. Note that the hot particle is never chosen in the sink regions.
Let $D$ the number of defects at the end of a mode. We will define the following Condition and use it in the proof of Theorem 1.
Condition 1.
@@
Condition 1 ensures that there is enough free thawed particles to keep the procedure running.
Proposition 2.
There is $c>0$ depending on $\lambda$ such that, if at the beginning of a mode Condition 1 is satisfied, then at the beginning of the next mode Condition 1 will also be satisfied with probability at least $1-e^{-cn}$, for all $n$ sufficiently large, uniformly on all possible configurations at the beginning of the mode.
We will prove Proposition 2 at the end of this section. Now we are ready to prove Theorem 1.
Proof of Theorem 1.
We take $1-\frac{1}{4K}<\zeta<1$. Since the initial configuration is deterministic with $\zeta N$ particles, at the initial configuration there are at most $n/4$ defects, and at least $3n/4$ blocks with one free thawed particle. Thus, there will be at least one block without defects and with one free thawed particle, and we can start the toppling procedure.
We will call a mode successful if Condition 1 is met at the end of the mode. We first note that, if a mode is successful, at least one jump is made. This happens because there is never more than one free thawed particle on a block which has defects. Thus, Condition 1 ensures that there are at least $n/4$ free particles on blocks without defects. At the end of the mode, such free thawed particles will be necessarily at the sink. Hence, at least one jump was made.
Using Proposition 2,
@@
Let $E_{n}$ be the set of blocks that realize at least one successful emission during a mode, $F(E_{n})$ be the number of frozen particles in $E_{n}$ and the end of a mode and $F(\mathbb{Z}_{N})$ the total number of frozen particles in the ring $\mathbb{Z}_{N}$ at the end of a mode.
To prove the Proposition 2 we will make use of the following statements.
Proposition 3.
There exists a constant $c$ such that, if Condition 1 is satisfied then $\mathbb{P}(F(E_{n})\geq\frac{n}{8})\leq e^{-cn}$ for all $n$ sufficiently large, uniformly on all possible configurations at the beginning of the mode.
Proposition 4.
There exists a constant $c>0$ such that, if at the beginning of the mode all the holes are at positions $iK$, for $i=1,\dots,n$, then $\mathbb{P}(F(\mathbb{Z}_{N})\geq\frac{n}{8})\leq e^{-cn}$ for all $n$ sufficiently large, uniformly on all possible configurations at the beginning of the mode.
Proof of Proposition 2.
Note that the number of free particles minus the number of defects is conserved by the toppling procedure, therefore at the end of the mode there are still at least $\frac{7}{8}n+D$ free particles.
In the first mode all the holes are at positions $iK$, for $i=1,\dots,n$, hence by Proposition 4, the probability of having at most $n/8$ frozen particles in total on the ring at the end of the mode is at least $1-e^{-cn}$. In this event, Condition 1 is met at the end of the mode.
Now we will consider the case from the second mode onwards. Let us show that, on the event $\{F(E_{n})\leq n/8\}$, Condition 1 is met at the end of the mode.
We first argue that in this event, there will be free particles inside the sink blocks. From the second mode onward, all free thawed particles begin the mode in the source blocks, except free thawed particles in blocks which still have defects, if any. Therefore, $E_{n}$ is a connected set surrounding the sources, and new frozen particles outside $E_{n}$ must be in the two blocks adjacent to $E_{n}$.
If $\{F(E_{n})\leq n/8\}$ happens, at the end of the mode, there are at most 2 new frozen particles outside $E_{n}$ and $5n/8$ particles that were frozen at the beginning of the mode. Since $5n/8+n/8+2<7n/8$ (if $n>16$), we conclude that the number of free thawed particles at the end of the mode is strictly higher than the number of defects. On a block with a defect there can be at most one free thawed particle; thus, the number of free thawed particles on blocks with defects is at most the number of defects, and there must have been free particles that are not frozen by the end of the mode, and this can only be achieved if free particles are emitted to a sink block.
Therefore, we conclude that, in fact, $E_{n}$ must contain the source block and be adjacent to a sink block, and since $E_{n}$ is a connected collection of blocks, the number of blocks in the complement of the union of $E_{n}$ and the sink blocks is at most $n/2$. In particular, by (P7), there are at most $n/2$ frozen particles outside $E_{n}$, hence at most $5n/8$ frozen particles in total.
By Proposition 3, $\mathbb{P}(F(E_{n})<n/8)\geq 1-e^{-cn}$, and Proposition 2 is proven. ∎